OpenStudy (anonymous):

Find the zeros for P(x) = 2x^3 - 7x^2 - 15x

OpenStudy (ash2326):

@odehye We have to find the solution of this \[2x^3-7x^2-15x=0\] Take x common \[x(2x^2-7x-15)=0\] Now we have \[x=0\ \text{or}\ (2x^2-7x-15)=0\] So one root is zero Can you solve this quadratic equation to find the other two roots ? \[2x^2-7x-15=0\]

OpenStudy (ash2326):

Do you get this part @odehye ?

OpenStudy (anonymous):

yes, i did

OpenStudy (ash2326):

Can you solve the quadratic to find the other two roots?

OpenStudy (anonymous):

i'm lost

OpenStudy (anonymous):

using P/Q theorem

OpenStudy (rob1525):

The easiest way is to us the quadratic formual to solve for zeros.

OpenStudy (anonymous):

x=5 and x=-3/2

OpenStudy (anonymous):

@ash2326 x=x=5 and x=-3/2